# Wen Wang    Ruibing Hou    Hong Chang    Shiguang Shan    Xilin Chen

### v2

@@ Abstract
Large audio-language models (LALMs), built upon powerful Large Language Models (LLMs), have exhibited remarkable audio comprehension and reasoning capabilities. However, the training of LALMs demands a large corpus of audio-language pairs, which requires substantial costs in both data collection and training resources. In this paper, we propose MATS, an audio-language multimodal LLM designed to handle Multiple Audio task using solely Text-only Supervision. By leveraging pre-trained audio-language alignment models such as CLAP, we develop a text-only training strategy that projects the shared audio-language latent space into LLM latent space, endowing the LLM with audio comprehension capabilities without relying on audio data during training. To further bridge the modality gap between audio and language embeddings within CLAP, we propose the Strongly-related noisy text with audio (Santa) mechanism. Santa maps audio embeddings into CLAP language embedding space while preserving essential information from the audio input. Extensive experiments demonstrate that MATS, despite being trained exclusively on text data, achieves competitive performance compared to recent LALMs trained on large-scale audio-language pairs.
@@ 1 Introduction
To reduce resource demands, researchers have proposed text-only audio captioning frameworks (Deshmukh et al., 2024; Kouzelis & Katsouros, 2023; Zhang et al., 2024; Li et al., 2024) which generate audio captions without relying on costly audio-language pairs for training. These frameworks leverage CLAP, which have established a well-aligned audio-language embedding space, enabling LLMs to interpret audio embedding by linking CLAP language embedding space with LLM semantic space. However, these methods have certain limitations. First, they are narrowly tailored to specific audio type (sound) and constrained to single audio captioning task, limiting their applicability to boarder audio open-ended question-answering scenarios. Second, to bridge the audio-language modality gap within CLAP, these methods either rely solely on noise injection during training or exclusively utilize a memory bank during inference. However, using only noise injection is highly random and sensitive, while relying solely on a memory bank during inference not only sacrifices audio embedding information but also increases inference time, ultimately compromising overall effectiveness.
@@ 2 Related work
Text-only Supervised Multimodal LLMs.  To alleviate resource demands, researchers have proposed zero-shot captioning frameworks, aimed at generating image/audio captions through text-only training. In the visual domain, large-scale pre-trained contrastive models like CLIP (Radford et al., 2021) align images and language into a shared vision-language embedding space. Building on CLIP, CapDec (Nukrai et al., 2022) trains a decoder to reconstruct text from its corresponding CLIP language embedding, which is then used to decode CLIP image embeddings at inference. To mitigate the vision-language modality gap, CapDec injects noise into language embedding during training. Differently, DeCap (Li et al., 2023) leverages a memory to store CLIP language embeddings, which is subsequently used to project visual embedding into CLIP language embedding space at inference. In the audio domain, models such as NoAudioCaptioning (Deshmukh et al., 2024), WSAC (Kouzelis & Katsouros, 2023), PromptACC (Zhang et al., 2024), and DRCap (Li et al., 2024) adopt similar strategies, replacing CLIP with CLAP to develop zero-shot audio captioning frameworks. However, these works focus on audio captioning, lacking the flexibility to handle a diverse range of audio tasks simultaneously, and they are limited to processing only a single type of audio (sound).
@@ 3.3 Theoretical Analysis on Generalization
[…]
@@ 3.4 Santa Mechanism
To obtain the CLAP language embedding space augmented with noise injection, we randomly select $M$ texts from the text-only training set, denoted as $\left\{T_{i}\right\}_{i=1}^{M}$, and construct a memory =$\mathcal{M}=\left\{\mathbf{m_{i}}\right\}_{i=1}^{M}$, where $\mathbf{m_{i}}=\mathcal{E}_{T}\left(T_{i}\right)+\epsilon$. At inference, the goal is to generate an answer for a given audio $A$ and associated instruction. Specifically, given the audio embedding $z^{a}=\mathcal{E}_{A}\left(A\right)$, we obtain its representation in language embedding space by combing $z^{a}$ with a weighted sum of semantically relevant embeddings in memory $\mathcal{M}$. A straightforward strategy to obtain the weights of these language embeddings is to compute their similarity with the audio embedding. However, due to the limited representational power of individual language embedding, this strategy is prone to retrieving the texts with insufficient semantic relevance, thereby affecting the effectiveness of audio-language modality alignment.
To this end, we apply k-means clustering to $\mathcal{M}$ and obtain a clustered memory $\mathcal{S}=\left\{\mathcal{S}_{k}\right\}_{k=1}^{K}$, where $\mathcal{S}_{k}$ stores the augmented language embeddings assigned to this cluster. By aggregating semantically similar augmented language embeddings, each cluster exhibits enhanced semantic representational capabilities. Denote the cluster centers as $\left\{\mathbf{c_{k}}\right\}_{k=1}^{K}$, we identity the most relevant cluster $\mathcal{S}_{\mathrm{closest}}$ by calculating the distance between the audio embedding $z^{a}$ and the cluster centers:
@@ 3.5 Training Pipeline
AudioTIA-5M Dataset.  MATS is designed to handle a wide range of audio tasks relying solely on text supervision, which makes existing audio datasets unsuitable for direct use in its training. In this work, we construct a new dataset AudioTIA-5M, where each training sample is formatted as (text, instruction, answer) tuple. AudioTIA-5M encompasses both close-ended and open-ended tasks, employing four audio-task templates in Table S8 of Appendix. For the classification task (CLS), we leverage ChatGPT (Achiam et al., 2023) to generate a substantial amount of textual data. Specifically, by utilizing ChatGPT’s instruction-following and text generation capabilities, we create a diverse set of audio descriptions (used as text) based on given event labels (used as answer). Further details are provided in Section B.4. For the audio captioning task (CAP), we directly use audio caption annotations from public datasets, including Clotho-v2 (Drossos et al., 2020), AudioCaps (Kim et al., 2019), MusicCaps (Agostinelli et al., 2023), WavCaps (Mei et al., 2024) and Macs (Martín-Morató & Mesaros, 2021), as text and answer. To further enhance model’s understanding and reasoning capabilities for audio, we incorporate two open-ended datasets: OpenAQA (Gong et al., 2024) and MusicQA (Liu et al., 2024a). In OpenAQA, we replace raw audio files with their audio captions as model inputs. For MusicQA, where music captions are unavailable, we generate corresponding captions using LP-MusicCaps (Doh et al., 2023). Overall, AudioTIA-5M comprises two subsets: a $1.5$ M close-ended question subset, and a $3.8$ M open-ended question subset. The detailed statistics of AudioTIA-5M are presented in Table S5 of Appendix.
@@
Table 1: Comparision results on audio classification benchmarks and ClothoAQA benchmark. $-{}^{\text{ZS-}}$ and indicate weakly zero-shot and supervised settings, respectively. Results without any annotation represent zero-shot setting. Note: the results of SALMONN and Qwen-Audio-Chat are from GAMA (Ghosh et al., 2024).
@@
[…]
@@ 4.2.1 Close-ended Results
Table 2: Comparision results for audio captioning tasks on AudioCaps and Clotho benchmarks.
@@
CIDEr SPICE SPICEr CIDEr SPICE SPICEr ROUGH-L BLUE4
@@
Audio Captioning.  Table 2 summarizes the comparative results for audio captioning tasks on AudioCaps, Clotho, and MusicCaps benchmarks. We can observe that: (1) Compared to text-only supervised audio captioning models, which are specifically tailored for audio captioning and restricted to the Sound type, our model demonstrates enhanced versatility by effectively addressing a broader spectrum of audio-relevant tasks across both Sound and Music type. Despite its broader applicability, our model achieves superior CIDEr and SPICEr scores on AudioCaps and a higher SPICE score on Clotho. The performance gains can be largely attributed to the proposed Santa mechanism, which effectively mitigates the audio-language modality gap within CLAP. (2) Compared to audio-supervised models, MATS-LLaMA exhibits comparable performance despite being training only on text data. Also, MATS-GPT2 surpasses the music captioning model LP-MusicCaps (Doh et al., 2023) by 9.9% on the ROUGH-L metric. These results underscore the effective audio captioning capabilities of MATS under text-only supervision.
@@ 4.2.2 Open-ended Results
To further evaluate our model’s understanding and reasoning capabilities in complex audio QA tasks, we conducted experiments on AIR-Bench Chat (Yang et al., 2024) and MMAU (Sakshi et al., 2024) benchmarks. The results are presented in Section 4.2.2. (1) On AIR-Bench, under text-only supervision, MATS-LLaMA achieves performance comparable to Qwen-Audio (Chu et al., 2023), ranking third-highest on Sound type and fourth-highest on Music type. We argue that CLAP struggles to extract fine-grained audio information, such as the exact number of occurrences. This limitation makes MATS-LLaMA more prone to errors when responding to fine-grained questions on AIR-Bench. (2) For the MMAU, the QA format requires LALMs to select from a set of provided options. We use the following prompt: {question} + Select one option: {options} Respond with exactly one of the options above. Compared to AIR-Bench, this option-based format offers additional contextual cues, partially mitigating CLAP’s limitations in extracting detailed information. Consequently, our model outperforms Qwen-Audio-Chat (Bai et al., 2023) by about $3\%$ in the sound type and ranks second only to the Qwen2-Audio-Instruction (Chu et al., 2024) in other audio types. Notably, our model even surpasses larger-scale models like Qwen-Audio-Chat and SALMONN. Furthermore, compared to MATS-Audio, MATS-LLaMA achieves comparable results on open-ended tasks, further
Table 3: Comparison results for open-ended audio tasks across sound and music types.
validating the effectiveness of the text-supervised strategy for training LALMs.
@@ 4.3 Ablation
MATS-GPT2 vs MATS-LLaMA.  As shown in Table 1 and Table 2 , MATS-GPT2 achieves performance comparable to MATS-LLaMA on close-ended tasks. However, as shown in Section 4.2.2, MATS-GPT2 exhibits significantly weak performance on open-ended tasks. These results indicate that a smaller LLM like GPT2 is sufficient for handling basic audio perception tasks, but it struggles with complex open-ended tasks that require advanced audio understanding and reasoning. In contrast, larger LLMs are better equipped to capture the depth and complexity necessary for audio reasoning and comprehension, making them indispensable for open-ended tasks.
@@
MATS-LLaMA vs MATS-Audio.  As shown in Table 1, MATS-LLaMA achieves performance comparable to MATS-Audio in the supervised setting. However, under the zero-shot setting, MATS-LLaMA significantly outperforms MATS-Audio, highlighting the heavy reliance of audio-supervised models on audio-language pairs for achieving superior performance. Moreover, as shown in Section 4.2.2, MATS-LLaMA achieves performance on par with MATS-Audio in open-ended scenarios, further highlighting the cost-effectiveness and efficiency of our approach.
@@
Santa vs other modal-transfer mechanism.  Section 4.3 presents the ablation study of our proposed Santa mechanism. The study evaluates the impact of excluding various components, including k-means (w/o K-means), memory bank (w/o MB), audio embedding (w/o AE) and Gaussian noise (w/o GN). As shown in Section 4.3, removing any component of Santa results in a performance drop, highlighting their collective effectiveness. Notably, the absence of Gaussian noise (w/o GN) during training and the exclusion of audio embeddings (w/o AE) during inference lead to significant performance decline. These results emphasize the critical role of incorporating Gaussian noise during training and retaining audio embeddings during inference to achieve robust performance. Additionally, we compare Santa with other modal-transfer methods, including the memory-based mechanism in DRCap (Li et al., 2024) and the noise-based mechanism in NoAudioCaptioning (Deshmukh et al., 2024). As shown in Section 4.3, Santa achieves better performance, validating its superiority in bridging the audio-language modality gap.
@@
The impact of Gaussian noise variance $\sigma$.  Section 4.3 investigates the impact of the variance $\sigma$ of injected noise during training. MATS achieves the best performance at $\sigma=0.015$. We argue that a small variance fails to effectively mitigate the audio-language modality gap within CLAP, while an excessively large variance impairs the model’s ability to capture useful input information.
@@
Figure 4: The impact of $\sigma$ on AudioCaps benchmark.

### v3

@@ Abstract
Large audio-language models (LALMs), built upon powerful Large Language Models (LLMs), have exhibited remarkable audio comprehension and reasoning capabilities. However, the training of LALMs demands a large corpus of audio-language pairs, which requires substantial costs in both data collection and training resources. In this paper, we propose MATS, an audio-language multimodal LLM designed to handle Multiple Audio task using solely Text-only Supervision. By leveraging pre-trained audio-language alignment models such as CLAP, we develop a text-only training strategy that projects the shared audio-language latent space into LLM latent space, endowing the LLM with audio comprehension capabilities without relying on audio data during training. To further bridge the modality gap between audio and language embeddings within CLAP, we propose the Strongly-related noisy text with audio (Santa) mechanism. Santa maps audio embeddings into CLAP language embedding space while preserving essential information from the audio input. Extensive experiments demonstrate that MATS, despite being trained exclusively on text data, achieves competitive performance compared to recent LALMs trained on large-scale audio-language pairs. The code is publicly available in https://github.com/wangwen-banban/MATS.
@@ 1 Introduction
To reduce resource demands, researchers have proposed text-only audio captioning frameworks (Deshmukh et al., 2024; Kouzelis and Katsouros, 2023; Zhang et al., 2024; Li et al., 2024) which generate audio captions without relying on costly audio-language pairs for training. These frameworks leverage CLAP, which have established a well-aligned audio-language embedding space, enabling LLMs to interpret audio embedding by linking CLAP language embedding space with LLM semantic space. However, these methods have certain limitations. First, they are narrowly tailored to specific audio type (sound) and constrained to single audio captioning task, limiting their applicability to boarder audio open-ended question-answering scenarios. Second, to bridge the audio-language modality gap within CLAP, these methods either rely solely on noise injection during training or exclusively utilize a memory bank during inference. However, using only noise injection is highly random and sensitive, while relying solely on a memory bank during inference not only sacrifices audio embedding information but also increases inference time, ultimately compromising overall effectiveness.
@@ 2 Related work
Text-only Supervised Multimodal LLMs. To alleviate resource demands, researchers have proposed zero-shot captioning frameworks, aimed at generating image/audio captions through text-only training. In the visual domain, large-scale pre-trained contrastive models like CLIP (Radford et al., 2021) align images and language into a shared vision-language embedding space. Building on CLIP, CapDec (Nukrai et al., 2022) trains a decoder to reconstruct text from its corresponding CLIP language embedding, which is then used to decode CLIP image embeddings at inference. To mitigate the vision-language modality gap, CapDec injects noise into language embedding during training. Differently, DeCap (Li et al., 2023) leverages a memory to store CLIP language embeddings, which is subsequently used to project visual embedding into CLIP language embedding space at inference. In the audio domain, models such as NoAudioCaptioning (Deshmukh et al., 2024), WSAC (Kouzelis and Katsouros, 2023), PromptAAC (Zhang et al., 2024), and DRCap (Li et al., 2024) adopt similar strategies, replacing CLIP with CLAP to develop zero-shot audio captioning frameworks. However, these works focus on audio captioning, lacking the flexibility to handle a diverse range of audio tasks simultaneously, and they are limited to processing only a single type of audio (sound).
@@ 3.3 Theoretical Analysis on Generalization
where $\left|\cdot\right|$ is $L_{1}$ norm. With these definitions, we can derive the generalization risk bound of the text-only supervised audio model, leading to following theorem.
@@ 3.4 Santa Mechanism
To obtain the CLAP language embedding space augmented with noise injection, we randomly select $M$ texts from the text-only training set, denoted as $\left\{T_{i}\right\}_{i=1}^{M}$, and construct a memory $\mathcal{M}=\left\{\mathbf{m_{i}}\right\}_{i=1}^{M}$, where $\mathbf{m_{i}}=\mathcal{E}_{T}\left(T_{i}\right)+\epsilon$. At inference, the goal is to generate an answer for a given audio $A$ and associated instruction. Specifically, given the audio embedding $z^{a}=\mathcal{E}_{A}\left(A\right)$, we obtain its representation in language embedding space by combing $z^{a}$ with a weighted sum of semantically relevant embeddings in memory $\mathcal{M}$. A straightforward strategy to obtain the weights of these language embeddings is to compute their similarity with the audio embedding. However, due to the limited representational power of individual language embedding, this strategy is prone to retrieving the texts with insufficient semantic relevance, thereby affecting the effectiveness of audio-language modality alignment.
To this end, we apply k-means clustering to $\mathcal{M}$ and obtain a clustered memory $\mathcal{S}=\left\{\mathcal{S}_{k}\right\}_{k=1}^{K}$, where $\mathcal{S}_{k}$ stores the augmented language embeddings assigned to this cluster. By aggregating semantically similar augmented language embeddings, each cluster exhibits enhanced semantic representational capabilities. Denote the cluster centers as $\left\{\mathbf{c_{k}}\right\}_{k=1}^{K}$, we identify the most relevant cluster $\mathcal{S}_{\mathrm{closest}}$ by calculating the distance between the audio embedding $z^{a}$ and the cluster centers:
@@ 3.5 Training Pipeline
AudioTIA-5M Dataset. MATS is designed to handle a wide range of audio tasks relying solely on text supervision, which makes existing audio datasets unsuitable for direct use in its training. In this work, we construct a new dataset AudioTIA-5M, where each training sample is formatted as (text, instruction, answer) tuple. AudioTIA-5M encompasses both close-ended and open-ended tasks, employing four audio-task templates in Table S13 of Appendix. For the classification task (CLS), we leverage ChatGPT (Achiam et al., 2023) to generate a substantial amount of textual data. Specifically, by utilizing ChatGPT’s instruction-following and text generation capabilities, we create a diverse set of audio descriptions (used as text) based on given event labels (used as answer). Further details are provided in Section B.4. For the audio captioning task (CAP), we directly use audio caption annotations from public datasets, including Clotho-v2 (Drossos et al., 2020), AudioCaps (Kim et al., 2019), MusicCaps (Agostinelli et al., 2023), WavCaps (Mei et al., 2024) and Macs (Martín-Morató and Mesaros, 2021), as text and answer. To further enhance model’s understanding and reasoning capabilities for audio, we incorporate two open-ended datasets: OpenAQA (Gong et al., 2024) and MusicQA (Liu et al., 2024a). In OpenAQA, we replace raw audio files with their audio captions as model inputs. For MusicQA, where music captions are unavailable, we generate corresponding captions using LP-MusicCaps (Doh et al., 2023). Overall, AudioTIA-5M comprises two subsets: a $1.5$ M close-ended question subset, and a $3.8$ M open-ended question subset. The detailed statistics of AudioTIA-5M are presented in Table S10 of Appendix.
@@
Table 1: Comparision results on audio classification benchmarks and ClothoAQA benchmark. $^{\text{ZS-}}$ and indicate weakly zero-shot and supervised settings, respectively. Results without any annotation represent zero-shot setting. Note: the results of SALMONN and Qwen-Audio-Chat are from GAMA (Ghosh et al., 2024).
@@
Model ESC-50 (ACC) FSD50K (mAP) DCASE (ACC) TUT (ACC) VGG (ACC) US8K (ACC) ClothoAQA ($\text{ACC}|\text{B-ACC}$) GTZAN (ACC) BJO (ACC)
@@ 4.2.1 Close-ended Results
Table 2: Comparision results for audio captioning tasks on AudioCaps, Clotho and MusicCaps benchmarks.
@@
CIDEr SPICE SPIDEr CIDEr SPICE SPIDEr ROUGH-L BLUE4
@@
Audio Captioning. Table 2 summarizes the comparative results for audio captioning tasks on AudioCaps, Clotho, and MusicCaps benchmarks. We can observe that: (1) Compared to text-only supervised audio captioning models, which are specifically tailored for audio captioning and restricted to the Sound type, our model demonstrates enhanced versatility by effectively addressing a broader spectrum of audio-relevant tasks across both Sound and Music type. Despite its broader applicability, our model achieves superior CIDEr and SPIDEr scores on AudioCaps and a higher SPICE score on Clotho. The performance gains can be largely attributed to the proposed Santa mechanism, which effectively mitigates the audio-language modality gap within CLAP. (2) Compared to audio-supervised models, MATS-LLaMA exhibits comparable performance despite being training only on text data. Also, MATS-GPT2 surpasses the music captioning model LP-MusicCaps (Doh et al., 2023) by 9.9% on the ROUGH-L metric. These results underscore the effective audio captioning capabilities of MATS under text-only supervision.
@@
Table 3: The results of our model on the MMAU datasets across the sound and music categories.
Method Size Sound Music Avg
Test-mini Test Test-mini Test Test-mini Test
AS Qwen2-Audio-Instruction (Chu et al., 2024) 8.4B 55.0 45.9 51.0 53.3 49.2 52.5
Qwen-Audio-Chat (Bai et al., 2023) 8.4B 55.3 56.7 44.0 40.9 43.1 41.9
SALMONN (Tang et al., 2024) 13B 41.0 40.3 34.8 33.8 33.7 32.8
GAMA (Ghosh et al., 2024) 7B 41.4 45.4 32.3 30.8 30.9 31.8
MuLLaMa (Liu et al., 2024b) 7B 40.8 44.8 32.6 30.6 31.9 30.7
GAMA-IT (Ghosh et al., 2024) 7B 43.2 43.2 28.4 28.0 30.2 29.0
LTU-AS (Gong et al., 2024) 7B 23.4 25.0 9.1 10.5 17.7 18.9
LTU (Gong et al., 2024) 7B 22.5 25.9 9.7 12.8 16.9 18.5
Pengi (Deshmukh et al., 2023) 323M 6.1 8.0 2.9 3.1 3.4 4.2
TS MATS-GPT2 (Ours) 370M 7.8 7.2 14.1 11.7 10.5 9.8
MATS-LLaMA (Ours) 7B 52.3 59.8 44.9 42.9 43.3 44.5
AS MATS-Audio 7B 51.7 57.4 43.7 41.5 42.2 43.8
Table 4: The results of our model on the MMAU datasets across the sound and music categories.
Method Size Sound Music Avg
Test-mini Test Test-mini Test Test-mini Test
AS Qwen2-Audio-Instruction 8.4B 55.0 45.9 51.0 53.3 49.2 52.5
Qwen-Audio-Chat 8.4B 55.3 56.7 44.0 40.9 43.1 41.9
SALMONN 13B 41.0 40.3 34.8 33.8 33.7 32.8
GAMA 7B 41.4 45.4 32.3 30.8 30.9 31.8
MuLLaMa 7B 40.8 44.8 32.6 30.6 31.9 30.7
GAMA-IT 7B 43.2 43.2 28.4 28.0 30.2 29.0
LTU-AS 7B 23.4 25.0 9.1 10.5 17.7 18.9
LTU 7B 22.5 25.9 9.7 12.8 16.9 18.5
Pengi 323M 6.1 8.0 2.9 3.1 3.4 4.2
TS MATS-GPT2 (Ours) 370M 7.8 7.2 14.1 11.7 10.5 9.8
MATS-LLaMA (Ours) 7B 52.3 59.8 44.9 42.9 43.3 44.5
AS MATS-Audio 7B 51.7 57.4 43.7 41.5 42.2 43.8
@@ 4.2.2 Open-ended Results
To further evaluate our model’s understanding and reasoning capabilities in complex audio QA tasks, we conducted experiments on AIR-Bench Chat (Yang et al., 2024) and MMAU (Sakshi et al., 2024) benchmarks. The results are presented in Table 5 and Table 3. (1) On AIR-Bench, under text-only supervision, MATS-LLaMA achieves performance comparable to Qwen-Audio (Chu et al., 2023), ranking third-highest on Sound type and fourth-highest on Music type. We argue that CLAP struggles to extract fine-grained audio information, such as the exact number of occurrences. This limitation makes MATS-LLaMA more prone to errors when responding to fine-grained questions on AIR-Bench. (2) For the MMAU, the QA format requires LALMs to select from a set of provided options. We use the following prompt: {question} + Select one option: {options} Respond with exactly one of the options above. Compared to AIR-Bench, this option-based format offers additional contextual cues, partially mitigating CLAP’s limitations in extracting detailed information. Consequently, our model outperforms Qwen-Audio-Chat (Bai et al., 2023) by about $3\%$ in the sound type and ranks second only to the Qwen2-Audio-Instruction (Chu et al., 2024) in other audio types. Notably, our model even surpasses larger-scale models like Qwen-Audio-Chat and SALMONN. Furthermore, compared to MATS-Audio, MATS-LLaMA achieves comparable results on open-ended tasks, further validating the effectiveness of the text-supervised strategy for training LALMs.
Table 5: The results of our model on the AIR-Bench Chat Benchmark across the sound and music categories.
Method Sound Music
AS Qwen2-Audio (Chu et al., 2024) 6.99 6.79
Qwen-Audio-Turbo (Bai et al., 2023) 6.59 5.98
SALMONN (Tang et al., 2024) 6.28 5.95
Qwen-Audio (Chu et al., 2023) 6.95 5.52
Gemini-1.5-pro (Team et al., 2024) 5.49 5.06
BLSP (Wang et al., 2023) 5.55 5.08
Pandagpt (Su et al., 2023) 5.46 5.06
Next-gpt (Wu et al., 2024) 4.76 4.18
SpeechGPT (Zhang et al., 2023) 0.95 0.95
Macaw-LLM (Lyu et al., 2023) 1.01 0.91
TS MATS-GPT2 (Ours) 4.33 3.31
MATS-LLaMA (Ours) 6.43 5.76
AS MATS-Audio 6.25 5.77
Table 6: The results of our model on the AIR-Bench Chat Benchmark across the sound and music categories.
Method Sound Music
AS Qwen2-Audio 6.99 6.79
Qwen-Audio-Turbo 6.59 5.98
SALMONN 6.28 5.95
Qwen-Audio 6.95 5.52
Gemini-1.5-pro 5.49 5.06
BLSP 5.55 5.08
Pandagpt 5.46 5.06
Next-gpt 4.76 4.18
SpeechGPT 0.95 0.95
Macaw-LLM 1.01 0.91
TS MATS-GPT2 (Ours) 4.33 3.31
MATS-LLaMA (Ours) 6.43 5.76
AS MATS-Audio 6.25 5.77
@@ 4.3 Ablation
MATS-GPT2 vs MATS-LLaMA. As shown in Table 1 and Table 2 , MATS-GPT2 achieves performance comparable to MATS-LLaMA on close-ended tasks. However, as shown in Table 5 and Table 3, MATS-GPT2 exhibits significantly weak performance on open-ended tasks. These results indicate that a smaller LLM like GPT2 is sufficient for handling basic audio perception tasks, but it struggles with complex open-ended tasks that require advanced audio understanding and reasoning. In contrast, larger LLMs are better equipped to capture the depth and complexity necessary for audio reasoning and comprehension, making them indispensable for open-ended tasks.
@@
MATS-LLaMA vs MATS-Audio. As shown in Table 1, MATS-LLaMA achieves performance comparable to MATS-Audio in the supervised setting. However, under the zero-shot setting, MATS-LLaMA significantly outperforms MATS-Audio, highlighting the heavy reliance of audio-supervised models on audio-language pairs for achieving superior performance. Moreover, as shown in Table 5 and Table 3, MATS-LLaMA achieves performance on par with MATS-Audio in open-ended scenarios, further highlighting the cost-effectiveness and efficiency of our approach.
@@
Santa. Table 8 presents the ablation study of our proposed Santa mechanism. The study evaluates the impact of excluding various components, including k-means (w/o K-means), memory bank (w/o MB), audio embedding (w/o AE) and Gaussian noise (w/o GN). As shown in Table 8, removing any component of Santa results in a performance drop, highlighting their collective effectiveness. Notably, the absence of Gaussian noise (w/o GN) during training and the exclusion of audio embeddings (w/o AE) during inference lead to significant performance decline. These results emphasize the critical role of incorporating Gaussian noise during training and retaining audio embeddings during inference to achieve robust performance.
Table 7: Performance Comparison between Santa and Other Modality Transfer Methods on Close-Ended and Open-Ended Benchmarks
Benchmark ESC-50 AudioCaps AIRBench-Sound AIRBench-Music
Metric ACC CIDEr GPT4 GPT4
MATS-NoAudioCaptioning 0.88 0.698 6.15 5.71
MATS-PromptAAC 0.77 0.593 6.07 5.28
MATS-DRCap 0.84 0.619 5.83 5.29
MATS-LLaMA 0.88 0.735 6.43 5.76
Santa vs other modal-transfer mechanism. To demonstrate the superior effectiveness of the Santa mechanism, we compare Santa with other modal-transfer methods, including the memory-based mechanism in DRCap (Li et al., 2024), the noise-based mechanism in NoAudioCaptioning (Deshmukh et al., 2024) and PromptAAC (Zhang et al., 2024). As shown in Table 7, Santa consistently outperforms the competing methods, validating its strength in bridging the audio-language modality gap. In DRCap, the original audio embeddings are completely discarded during inference, making its performance highly dependent on the alignment between the memory bank and the test benchmark. However, in a multi-task setting, the memory bank must accommodate a diverse range of benchmarks rather than a single one, which introduces noise into the retrieval process and projects the audio embeddings into less relevant textual spaces, leading to performance drop. NoAudioCaptioning, on the other hand, employs a pure noise-injection strategy. While simple, this approach fails to effectively reduce the modality gap within CLAP, as analyzed in detail in Section C.1, resulting in limited performance gains. PromptAAC adopts an augmentation-based approach that involves injecting noise and substituting similar language inputs. It retrieves audio events by matching audio embeddings with language embeddings derived from 527 predefined audio labels in AudioSet. However, the limited variety of audio events restricts the diversity of the retrieved information, resulting in inferior performance compared to Santa.
@@
The impact of Gaussian noise variance $\sigma$. Table 9 investigates the impact of the variance $\sigma$ of injected noise during training. MATS achieves the best performance at $\sigma=0.015$. We argue that a small variance fails to effectively mitigate the audio-language modality gap within CLAP, while an excessively large variance impairs the model’s ability to capture useful input information.
@@
Table 9: The impact of $\sigma$ on AudioCaps benchmark.